# Real-Time Stochastic Terrain Mapping and Processing for Autonomous Safe Landing

### v1

@@ 4.3 Computational Cost Analysis
Table 7: Run time comparison between the proposed and baseline safety map generation algorithms and the proposed DEM generation algorithm across different map resolutions. The baseline DEM generation algorithm cannot specify the resolution. Thus, its run-time is not given.
Resolution Baseline (s) Proposed (s)
(m/pix) DEM Safety DEM Safety
@@
Even ignoring the DEM construction run time of the baseline algorithm, the proposed algorithm achieves faster computation except for the coarse resolution of 0.5 m/pix. In higher resolution cases, the sum of the run times of PCD to DEM and DEM to safety map by the proposed algorithm is smaller than the run time of DEM to safety map by the baseline algorithm. Comparing DEM to safety map run times between the resolution of 0.2 and 0.1 m/pix cases, we can confirm that the baseline algorithm takes $O(p^{5})$ while the proposed algorithm takes $O(p^{4})$, which aligns with the algorithm’s design as explained previously. Here, $p$ is the resolution, which is the number of pixels that span a unit length.
@@ 5 Conclusion
Appendix: Proof of Theorem 3.1

### v2

@@ 4.4 Computational Cost Analysis
Table 8: Run time comparison across different map resolutions.
Resolution, Baseline, s $\downarrow$ Proposed, s $\downarrow$
m/pix DEM Safety DEM Safety
@@
  • The baseline DEM generation algorithm cannot specify the resolution.
@@
Even ignoring the DEM construction run time of the baseline algorithm, the proposed algorithm achieves faster computation except for the coarse resolution of 0.5 m/pix. In higher resolution cases, the sum of the run times of PCD to DEM and DEM to safety map by the proposed algorithm is smaller than the run time of DEM to safety map by the baseline algorithm. Comparing DEM to safety map run times between the resolution of 0.2 and 0.1 m/pix cases, we can confirm that the baseline algorithm takes $O(p^{5})$ while the proposed algorithm takes $O(p^{4})$, which aligns with the algorithm’s design (Appendix 5.1). Here, $p$ is the resolution, which is the number of pixels that span a unit length.
@@ 5 Conclusion
Appendix
5.1 Scalability of Hazard Detection Algorithms
For fixed sizes of the digital elevation map (DEM) length and the lander diameter, their pixel span scales proportionally to $p$, the unit length of the pixel size. To evaluate the roughness safety, we need to compute all the pixels within the lander footprint, which scales $O(p^{2})$. We repeat this process for all possible lander orientations and for every DEM pixel. The number of lander orientation steps is driven by the width of the lander in pixels. Therefore, the exact hazard detection (HD) algorithm scales as $O(p^{5})$ [48]. The proposed provably conservative HD algorithm allows us to conservatively skip repeated evaluation for all possible orientations, resulting in scalability $O(p^{4})$.
5.2 Proof of Theorem 3.1
@@ 5.1 Scalability of Hazard Detection Algorithms
5.3 Realistic Terrain Data
To evaluate the algorithm’s performance on realistic data, we selected ten distinct locations on the Mars Digital Terrain Model (DTM) for the candidate ExoMars landing site in Hypanis Valles [62], as illustrated in Fig. 27. Randomly placed rocks were superimposed on these terrains to simulate realistic surface conditions. Figures 28 and 29 present the corresponding DEMs and safety maps for the selected regions.
Figure 27: Ten regions selected for high-fidelity analysis.
Figure 28: Digital Elevation Models (DEMs) of the selected regions.
Figure 29: Safety maps corresponding to the DEMs of the selected regions where white and black indicate safe and unsafe locations, respectively.